# Generalized Moving Least-Squares for Solving Vector-valued PDEs on Unknown Manifolds

### v1

@@ 4.3 Reduction from ambient dimension to intrinsic dimension
Using all these block submatrices $\hat{\mathbf{R}}_{sr}^{\ell}$ for $1\leq s,r\leq K$ and $\ell=1,2,3$, one can find the weights $\{\bm{w}^{k}\in\mathbb{R}^{2\times 2}\}_{k=1}^{K}$ satisfying $\Delta_{B}u({\mathbf{x}_{0}})\approx\sum_{k=1}^{K}\bm{w}^{k}u({\mathbf{x}_{0,k}})$ in (12) by
@@
The complete procedure of the extrinsic GMLS approximation of the Bochner Laplacian is listed in Algorithm 2. See Section B.2 and Section B.3 for the extrinsic approximations of two other vector Laplacians acting on vector fields.
@@
1:  Input: A point cloud $\{\mathbf{x}_{i}\}_{i=1}^{N}\subset M$, bases of (estimated) tangent vector space at each node $\left\{\bm{t}_{1}(\mathbf{x}_{i}),\bm{t}_{2}(\mathbf{x}_{i})\right\}_{i=1}^{N}$, (estimated) projection matrices $\mathbf{P}$, the degree $l$ of polynomials, and a parameter $K>(l+2)(l+1)/2$ nearest neighbors.
2:  Set $\mathbf{L}_{B}$ to be a sparse $2N\times 2N$ matrix with $4NK$ nonzeros.
@@
5:     Construct the differential matrix $\mathbf{G}_{\ell}$ as in (27).
6:     Calculate the block submatrices $\mathbf{\hat{R}}^{\ell}_{sr}$ as in (33) for $31\leq s,r\leq K,\ell=1,2,$ 3.
7:     Find the weights $\{\bm{w}^{k}\}_{k=1}^{K}$ as in (34).
@@ 4.4 Approximation of covariant derivative
We now compute the covariant derivative of a vector field $u$ along itself at the base point $\mathbf{x}_{0}$, $\nabla_{u}u(\mathbf{x}_{0})$, in extrinsic formulation using the $K$ neighboring points of $\mathbf{x}_{0}$ in $S_{\mathbf{x}_{0}}$. Let $U=(U^{1},U^{2},U^{3})^{\top}$ be an extension of $u=u^{j}\bm{t}_{j}$ so that $(U^{1},U^{2},U^{3})^{\top}=u^{1}\bm{t}_{1}+u^{2}\bm{t}_{2}$. The extrinsic formulation for the covariant derivative can be written as (see e.g. [22]),
@@
$\nabla_{u}u(\mathbf{x}_{0})=\mathbf{P}\left(\bar{\nabla}_{U}U\right)|_{\mathbf{x}_{0}}=\mathbf{TT}^{\top}\left.\left[\begin{matrix}\frac{\partial U^{1}}{\partial X^{1}}&\frac{\partial U^{1}}{\partial X^{2}}&\frac{\partial U^{1}}{\partial X^{3}}\\ \frac{\partial U^{2}}{\partial X^{1}}&\frac{\partial U^{2}}{\partial X^{2}}&\frac{\partial U^{2}}{\partial X^{3}}\\ \frac{\partial U^{3}}{\partial X^{1}}&\frac{\partial U^{3}}{\partial X^{2}}&\frac{\partial U^{3}}{\partial X^{3}}\end{matrix}\right]\left[\begin{matrix}U^{1}\\ U^{2}\\ U^{3}\end{matrix}\right]\right|_{\mathbf{x}_{0}},$ (35)
@@
where $\mathbf{T}=[\bm{t}_{1},\bm{t}_{2}]\in\mathbb{R}^{3\times 2}$ is the global tangent vector basis. Then we approximate $\left[\frac{\partial U^{r}}{\partial X^{s}}\right]_{r,s=1}^{3}$ using GMLS approach. Notice that $\frac{\partial}{\partial X^{s}}=\left(\mathbf{e}_{s}\cdot\mathbf{I}\right)\cdot\overline{\mathrm{grad}}_{\mathbb{R}^{3}}$ can be obtained from $\mathcal{G}_{s}=\left(\mathbf{e}_{s}\cdot\mathbf{P}\right)\cdot\overline{\mathrm{grad}}_{\mathbb{R}^{3}}$ in (22) by replacing the projection $\mathbf{P}$ with the identity $\mathbf{I}$. Following the similar formula as in (27), one can construct a $K$ by $K$ differential matrix $\mathbf{D}_{s}$ to approximate $\frac{\partial}{\partial X^{s}}$ in $S_{\mathbf{x}_{0}}$. Taking the first row of $\mathbf{D}_{s}$, denoted by $\mathbf{D}_{s1}\in\mathbb{R}^{1\times K}$, one can approximate the value of $\frac{\partial U^{r}}{\partial X^{s}}|_{\mathbf{x}_{0}}$ as $\mathbf{D}_{s1}\mathbf{U}_{{\mathbf{x}_{0}}}^{r}$ for $1\leq r,s\leq 3$, where $\mathbf{U}_{{\mathbf{x}_{0}}}^{r}=(U^{r}({\mathbf{x}_{0,1}}),...,U^{r}(\mathbf{x}_{0,K}))^{\top}\in\mathbb{R}^{K\times 1}$ has been defined right after (30). Then the covariant derivative (35) can be calculated as
$\nabla_{u}u(\mathbf{x}_{0})$ $=$ $\mathbf{T(\mathbf{x}_{0})T}(\mathbf{x}_{0})^{\top}\left[\begin{matrix}\mathbf{D}_{11}\mathbf{U}_{{\mathbf{x}_{0}}}^{1}&\mathbf{D}_{21}\mathbf{U}_{{\mathbf{x}_{0}}}^{1}&\mathbf{D}_{31}\mathbf{U}_{{\mathbf{x}_{0}}}^{1}\\ \mathbf{D}_{11}\mathbf{U}_{{\mathbf{x}_{0}}}^{2}&\mathbf{D}_{21}\mathbf{U}_{{\mathbf{x}_{0}}}^{2}&\mathbf{D}_{31}\mathbf{U}_{{\mathbf{x}_{0}}}^{2}\\ \mathbf{D}_{11}\mathbf{U}_{{\mathbf{x}_{0}}}^{3}&\mathbf{D}_{21}\mathbf{U}_{{\mathbf{x}_{0}}}^{3}&\mathbf{D}_{31}\mathbf{U}_{{\mathbf{x}_{0}}}^{3}\end{matrix}\right]\left[\begin{matrix}U^{1}(\mathbf{x}_{0})\\ U^{2}(\mathbf{x}_{0})\\ U^{3}(\mathbf{x}_{0})\end{matrix}\right]$
$=$ $\mathbf{T(\mathbf{x}_{0})T}(\mathbf{x}_{0})^{\top}\left[\begin{matrix}\left(\mathbf{U}_{{\mathbf{x}_{0}}}^{1}\right)^{\top}\\ \left(\mathbf{U}_{{\mathbf{x}_{0}}}^{2}\right)^{\top}\\ \left(\mathbf{U}_{{\mathbf{x}_{0}}}^{3}\right)^{\top}\end{matrix}\right]_{3\times K}\left[\begin{matrix}\mathbf{D}_{11}^{\top}&\mathbf{D}_{21}^{\top}&\mathbf{D}_{31}^{\top}\end{matrix}\right]_{K\times 3}\left[\begin{matrix}U^{1}(\mathbf{x}_{0})\\ U^{2}(\mathbf{x}_{0})\\ U^{3}(\mathbf{x}_{0})\end{matrix}\right].$
@@ 5 Numerical results
In this section, we show supporting numerical examples, including eigenvalue problems, Poisson equations, linear vector diffusion equations and nonlinear Burgers’ equations. In Section 5.1 we numerically investigate the eigenvalue stability of the Bochner Laplacian on manifolds. In Section 5.2, we verify the numerical convergence of the Bochner Laplacian for solving Poisson equations. In Section 5.3 and Section 5.4, we study the numerical performance of the linear vector diffusion equation and the nonlinear viscous Burgers’ equation, respectively.
In our numerical implementation, we only consider the regime of unknown manifolds, that is, we are only given a point cloud on unknown smooth manifolds. We apply the algorithm outlined in Section 2.2 to approximate the tangent space of the manifolds without alignment as shown in Fig. 1. For time-independent problems, we choose the polynomial degree of manifolds in GMLS to be identical to that used for functions. For time-dependent problems, the degree of manifolds is chosen to be 6 which is higher than that used for functions ($l=2,3,4,5$) such that the error from manifold approximation is negligible. Besides, the parameter $K$ in $K$-nearest neighbors are taken large enough, typically around $51\sim 71$, to yield a stable approximation of the Bochner Laplacian. The following numerical experiments demonstrate that the GMLS approaches can provide stable estimators and convergent results with small error bars, given various trials of independent randomly sampled data points.
@@ 5.1 Eigenvalue stability
In this subsection, we investigate the eigenvalue problem of the Bochner Laplacian on two examples. For PDE applications, a necessary condition for stability is that all eigenvalues of the discrete vector Laplacian have negative real parts. The first example is a 1D ellipse in $\mathbb{R}^{2}$ with the parameterization, $\mathbf{x}=(\cos\theta,2\sin\theta)$, where $0\leq\theta<2\pi$. For a 1D manifold, the Bochner Laplacian shares the same eigenvalues as the Hodge Laplacian, as well as the Laplace-Beltrami (see e.g., [22]). Here, the true eigenvalues can be approximated semi-analytically by solving the Sturm-Liouville form of the Laplace-Beltrami operator using the finite difference method (see p. 68 of [44]). The second example is a 2D unit sphere embedded in $\mathbb{R}^{3}$. We notice that the eigenvalues of the Bochner Laplacian are different from those of the Hodge Laplacian by a constant Ricci curvature on a unit sphere, where the non-trivial eigenvalues of the Hodge Laplacian are identical with the non-trivial eigenvalues of the Laplace-Beltrami operator but with double multiplicities (see e.g., [22]). In our numerical implementation, data points $\{\mathbf{x}_{i}\}_{i=1}^{N}$ are randomly sampled with a uniform distribution in intrinsic coordinates on these two manifolds.
In the first column of Fig. 2, we plot all the eigenvalues of the non-symmetric Bochner Laplacian matrices in the complex plane. We see that all the eigenvalues lie in the left half complex plane with non-positive real parts to achieve the eigenvalue stability of the GMLS approximation. In the second column of Fig. 2, we further observe that the leading 8 numerical eigenvalues are in good agreement with the semi-analytic true eigenvalues. However, when there is a zero eigenvalue, the GMLS approximation possibly provides only one corresponding value which is slightly greater than zero and all the other eigenvalues having strictly negative real parts (see $\hat{\lambda}_{0}=1.6\mathrm{e}-8$ in the ellipse example, where $\hat{\lambda}_{0}$ is a numerical eigenvalue with a positive real part close to zero.) In the following PDE applications, we always first test the eigenvalue stability before solving the PDE problems. If there is a numerical eigenvalue having small positive real part as in the ellipse example, we will regularize the Bochner Laplacian matrix by subtracting a constant from its diagonal entries, $\mathbf{L}_{B}-\hat{\lambda}_{0}\mathbf{I}$, to guarantee the eigenvalue stability where $\mathbf{I}$ is an identity matrix.
Ellipse
Unit Sphere
Ellipse
k True Extrinsic Intrinsic
1 0 1.6383e-08 -4.3835e-07
2 -0.4206 -0.4206 -0.4206
3 -0.4206 -0.4206 -0.4206
4 -1.6823 -1.6823 -1.6823
5 -1.6823 -1.6823 -1.6823
6 -3.7852 -3.7852 -3.7852
7 -3.7852 -3.7852 -3.7852
8 -6.7298 -6.7298 -6.7293
Unit Sphere
@@
Figure 2: Eigenvalues on 1D ellipse in $\mathbb{R}^{2}$ and 2D sphere in $\mathbb{R}^{3}$. In both examples, $K=51$, $l=5$, and $N=6400$ points are randomly sampled. The first column: all eigenvalues lie in the left half complex plane. The second column: comparison of the leading 8 eigenvalues of the Bochner Laplacian between our approximation using intrinsic and extrinsic methods and the semi-analytic true solution.
5.2 Poisson problems

### v2

@@ 4.3 Reduction from ambient dimension to intrinsic dimension
Using all these block submatrices $\hat{\mathbf{R}}_{sr}^{\ell}$ for $1\leq s,r\leq K$ and $\ell=1,2,3$, one can find the weights $\{\bm{w}^{k}\in\mathbb{R}^{2\times 2}\}_{k=1}^{K}$ satisfying $\Delta_{B}\bm{u}({\mathbf{x}_{0}})\approx\sum_{k=1}^{K}\bm{w}^{k}\bm{u}|_{{\mathbf{x}_{0,k}}}$ in (13) by
@@
See Section B.2 and Section B.3 for the extrinsic approximations of two other vector Laplacians acting on vector fields for 2D manifolds in $\mathbb{R}^{3}$.
4.4 Extrinsic formulation in higher dimensional spaces
Without loss of generality, the above extrinsic formulation can be readily extended to $d$-dimensional manifolds in $\mathbb{R}^{n}$. For simplicity, we present the generalized formulation of the Bochner Laplacian rather than other vector Laplacians. The complete procedure for the extrinsic GMLS approximation of the Bochner Laplacian as well as the complexity is outlined in Algorithm 2.
@@ 4.4 Extrinsic formulation in higher dimensional spaces
1:  Input: A point cloud $\{\mathbf{x}_{i}\}_{i=1}^{N}\subset M$, bases of (estimated) tangent vectors at each node $\left\{\bm{t}_{1}(\mathbf{x}_{i}),\ldots,\bm{t}_{d}(\mathbf{x}_{i})\right\}_{i=1}^{N}$, (estimated) projection matrices $\mathbf{P}$, the degree $l$ of polynomials, and a parameter $K(>m)$ nearest neighbors.
2:  Set $\mathbf{L}_{B}$ to be a sparse $dN\times dN$ matrix with $d^{2}NK$ nonzeros.
@@
5:     Construct the differential matrix $\mathbf{G}_{\ell}$ as in (39) with complexity $O(NnK^{2}m+NnKmd^{2})$.
6:     Calculate the block submatrices $\mathbf{\hat{R}}^{\ell}_{sr}\in\mathbb{R}^{d\times d}$ similarly as in (45) for $1\leq s,r\leq K$ and $\ell=1,\ldots,n$ with complexity $O(Nnd^{2}K^{2})$.
7:     Find the weights $\{\bm{w}^{k}\}_{k=1}^{K}$ similarly as in (46) with complexity $O(Nnd^{3}K^{2})$.
@@ 4.5 Approximation of covariant derivative
We now compute the covariant derivative of a vector field $\bm{u}$ along itself at the base point $\mathbf{x}_{0}$, $\nabla_{\bm{u}}\bm{u}(\mathbf{x}_{0})$, in extrinsic formulation using the $K$ neighboring points of $\mathbf{x}_{0}$ in $S_{\mathbf{x}_{0}}$. Let $\bm{U}=(U^{1},U^{2},U^{3})^{\top}$ be an extension of $\bm{u}=u^{j}\bm{t}_{j}$ so that $(U^{1},U^{2},U^{3})^{\top}=u^{1}\bm{t}_{1}+u^{2}\bm{t}_{2}$. The extrinsic formulation for the covariant derivative can be written as (see e.g. [24]),
@@
$\nabla_{\bm{u}}\bm{u}(\mathbf{x}_{0})=\mathbf{P}\left(\bar{\nabla}_{\bm{U}}\bm{U}\right)|_{\mathbf{x}_{0}}=\mathbf{TT}^{\top}\left.\left[\begin{matrix}\frac{\partial U^{1}}{\partial X^{1}}&\frac{\partial U^{1}}{\partial X^{2}}&\frac{\partial U^{1}}{\partial X^{3}}\\ \frac{\partial U^{2}}{\partial X^{1}}&\frac{\partial U^{2}}{\partial X^{2}}&\frac{\partial U^{2}}{\partial X^{3}}\\ \frac{\partial U^{3}}{\partial X^{1}}&\frac{\partial U^{3}}{\partial X^{2}}&\frac{\partial U^{3}}{\partial X^{3}}\end{matrix}\right]\left[\begin{matrix}U^{1}\\ U^{2}\\ U^{3}\end{matrix}\right]\right|_{\mathbf{x}_{0}},$ (47)
@@
where $\mathbf{T}=[\bm{t}_{1},\bm{t}_{2}]\in\mathbb{R}^{3\times 2}$ is the global tangent vector basis. Then we approximate $\left[\frac{\partial U^{r}}{\partial X^{s}}\right]_{r,s=1}^{3}$ using GMLS approach. Note that $\frac{\partial}{\partial X^{s}}=\left(\mathbf{e}_{s}\cdot\mathbf{I}\right)\cdot\overline{\mathrm{grad}}_{\mathbb{R}^{3}}$ can be obtained from $\mathcal{G}_{s}=\left(\mathbf{e}_{s}\cdot\mathbf{P}\right)\cdot\overline{\mathrm{grad}}_{\mathbb{R}^{3}}$ in (34) by replacing the projection $\mathbf{P}$ with the identity $\mathbf{I}$. Following the similar formula as in (39), one can construct a $K$ by $K$ differential matrix $\mathbf{D}_{s}$ to approximate $\frac{\partial}{\partial X^{s}}$ in $S_{\mathbf{x}_{0}}$. Taking the first row of $\mathbf{D}_{s}$, denoted by $\mathbf{D}_{s,1}\in\mathbb{R}^{1\times K}$, one can approximate the value of $\frac{\partial U^{r}}{\partial X^{s}}|_{\mathbf{x}_{0}}$ at the base $\mathbf{x}_{0}$ as $\mathbf{D}_{s,1}\mathbf{U}_{{\mathbf{x}_{0}}}^{r}$ for $1\leq r,s\leq 3$, where $\mathbf{U}_{{\mathbf{x}_{0}}}^{r}=(U^{r}({\mathbf{x}_{0,1}}),...,U^{r}(\mathbf{x}_{0,K}))^{\top}\in\mathbb{R}^{K\times 1}$ has been defined right after (42). Then the covariant derivative (47) can be calculated as
$\nabla_{\bm{u}}\bm{u}(\mathbf{x}_{0})$ $=$ $\mathbf{T(\mathbf{x}_{0})T}(\mathbf{x}_{0})^{\top}\left[\begin{matrix}\mathbf{D}_{1,1}\mathbf{U}_{{\mathbf{x}_{0}}}^{1}&\mathbf{D}_{2,1}\mathbf{U}_{{\mathbf{x}_{0}}}^{1}&\mathbf{D}_{3,1}\mathbf{U}_{{\mathbf{x}_{0}}}^{1}\\ \mathbf{D}_{1,1}\mathbf{U}_{{\mathbf{x}_{0}}}^{2}&\mathbf{D}_{2,1}\mathbf{U}_{{\mathbf{x}_{0}}}^{2}&\mathbf{D}_{3,1}\mathbf{U}_{{\mathbf{x}_{0}}}^{2}\\ \mathbf{D}_{1,1}\mathbf{U}_{{\mathbf{x}_{0}}}^{3}&\mathbf{D}_{2,1}\mathbf{U}_{{\mathbf{x}_{0}}}^{3}&\mathbf{D}_{3,1}\mathbf{U}_{{\mathbf{x}_{0}}}^{3}\end{matrix}\right]\left[\begin{matrix}U^{1}(\mathbf{x}_{0})\\ U^{2}(\mathbf{x}_{0})\\ U^{3}(\mathbf{x}_{0})\end{matrix}\right]$
$=$ $\mathbf{T(\mathbf{x}_{0})T}(\mathbf{x}_{0})^{\top}\left[\begin{matrix}\left(\mathbf{U}_{{\mathbf{x}_{0}}}^{1}\right)^{\top}\\ \left(\mathbf{U}_{{\mathbf{x}_{0}}}^{2}\right)^{\top}\\ \left(\mathbf{U}_{{\mathbf{x}_{0}}}^{3}\right)^{\top}\end{matrix}\right]_{3\times K}\left[\begin{matrix}\mathbf{D}_{1,1}^{\top}&\mathbf{D}_{2,1}^{\top}&\mathbf{D}_{3,1}^{\top}\end{matrix}\right]_{K\times 3}\left[\begin{matrix}U^{1}(\mathbf{x}_{0})\\ U^{2}(\mathbf{x}_{0})\\ U^{3}(\mathbf{x}_{0})\end{matrix}\right].$
@@ 5 Numerical results
To support the performance of both intrinsic and extrinsic GMLS approximations as well as to validate the error estimates in Remark 2.1, we present numerical examples, including eigenvalue problems, screened Poisson equations, linear vector diffusion equations, and nonlinear Burgers’ equations. In Section 5.1, we numerically investigate the eigenvalue stability of the vector Laplacian on manifolds. In Section 5.2, we verify the numerical convergence of the GMLS approximation for solving the vector screened Poisson equation formulated with the Bochner Laplacian and the Hodge Laplacian. In Section 5.3 and Section 5.4, we study the numerical performance of the linear vector diffusion equation and the nonlinear viscous Burgers’ equation, respectively.
In our numerical implementation, for known smooth manifolds, we are given a point cloud associated with its parametrization and tangent bundle. For unknown smooth manifolds, we are only given a point cloud and we apply the algorithm outlined in Section 2.2 to approximate the tangent vectors of the manifolds without alignment as shown in Fig. 2. Then we apply Algorithm 1 and Algorithm 2 to obtain intrinsic and extrinsic approximations of vector Laplacians, respectively. For time-independent problems, we choose the polynomial degree of manifolds in GMLS to be identical to that used for functions. For time-dependent problems, the degree of manifolds is chosen to be 6 which is higher than that used for functions ($l=2,3,4,5$) such that the error from manifold approximation is negligible. To achieve a stable and efficient approximation of the vector Laplacian, we empirically choose the $K$-nearest neighbors around $50\sim 70$ for 2D manifolds, satisfying $K>(l+2)(l+1)/2$ where degree $l=5$. The following numerical experiments demonstrate that the GMLS approaches can provide stable estimators and convergent results with small error bars, given various trials of independent randomly sampled data points.
@@ 5.1 Eigenvalue stability
Unit Sphere (Bochner)
Unit Sphere (Bochner)
@@
Figure 3: Eigenvalues on 2D sphere in $\mathbb{R}^{3}$. $K=50$, $l=5$, and $N=6400$ points are randomly sampled. (Left) All eigenvalues lie in the left half complex plane. (Right) Comparison of the leading 8 eigenvalues of the Bochner Laplacian between our approximation using intrinsic and extrinsic GMLS methods and the analytic truth.
In this subsection, we investigate the eigenvalue problem for vector Laplacians using two examples. For PDE applications, a necessary condition for stability is that all eigenvalues of the discrete vector Laplacian have non-positive real parts. The first example involves the Bochner Laplacian on a 2D unit sphere embedded in $\mathbb{R}^{3}$. The true eigenvalues of the Bochner Laplacian differ from those of the Hodge Laplacian by a constant Ricci curvature on the unit sphere. The non-trivial eigenvalues of the Hodge Laplacian coincide with those of the Laplace-Beltrami operator but with double multiplicities (see e.g., [24]). Neither of the two vector Laplacians has zero eigenvalues. In our numerical implementation, data points $\{\mathbf{x}_{i}\}_{i=1}^{N}$ are randomly sampled with a uniform distribution on the unit sphere. In the left panel of Fig. 3, we plot all the eigenvalues of the non-symmetric Bochner Laplacian matrices in the complex plane. We see that all the eigenvalues lie in the left half complex plane with negative real parts, ensuring the eigenvalue stability of the GMLS approximation. In the right panel of Fig. 3, we observe that the leading 8 numerical eigenvalues are in good agreement with the analytic true eigenvalues.
(a) Torus in $\mathbb{R}^{3}$ (Hodge)
Torus in $\mathbb{R}^{3}$ (Hodge)
k True Extrinsic Intrinsic
1 0 -2.7499e-04 -7.2051e-04
2 0 0.0010 -0.0050
3 -0.2494 -0.2480 -0.2504
4 -0.2494 -0.2481 -0.2504
5 -0.2494 -0.2494 -0.2549
6 -0.2494 -0.2495 -0.2551
7 -0.7946 -0.7930 -0.7955
8 -0.7946 -0.7931 -0.7956
(b) Two Leading Eigenvalues (c) 1st Eigenvector Field (d) 2nd Eigenvector Field
Refer to caption Refer to caption Refer to caption
Figure 4: Eigenvalues on 2D torus in $\mathbb{R}^{3}$. $K=50$, $l=5$, and $N=6400$ points are randomly sampled. The first row: (a) Almost all eigenvalues lie in the left half complex plane. (Right) Comparison of the leading 8 eigenvalues of the Hodge Laplacian between the GMLS and the semi-analytic truth. The second row: (b) The first two eigenvalues of our approximation (with one positive and the other negative) converge to zero with $O(N^{-2})$. (c) and (d) are the two corresponding harmonic fields of our approximation.
The second example involves the Hodge Laplacian on a torus parametrized by
$\mathbf{x}=\left[(R+r\cos\theta)\cos\phi,(R+r\cos\theta)\sin\phi,r\sin\theta\right],\quad 0\leq\theta<2\pi,0\leq\phi<2\pi,$ (48)
where $R$ is the distance from the center of the tube to the center of the torus and $r<R$ is the distance from the center of the tube to the surface of the tube. We set $R=2$ and $r=1$. The eigenvalues of the Hodge Laplacian are identical to those of the Laplace-Beltrami but with double multiplicities. Notably, the Hodge Laplacian on a torus has a zero eigenvalue with multiplicity two, associated with two harmonic fields. The nontrivial true eigenvalues can be approximated semi-analytically by solving the Sturm-Liouville form of the Laplace-Beltrami operator using the finite difference method (see p. 68 of [48]).
Numerically, the data $\{\theta_{i},\phi_{i}\}_{i=1}^{N}$ are generated from a uniform distribution on $\left[0,2\pi\right)\times\left[0,2\pi\right)$ and then mapped onto the point cloud $\{\mathbf{x}_{i}\}_{i=1}^{N}$ via the parametrization (48). We observe from the first row of Fig. 4 that almost all eigenvalues lie in the left half plane except a value of 0.001 which approximates the zero eigenvalue. Next, we study the behavior of the two numerical eigenvalues close to zero. From the second row of Fig. 4, we see that these two numerical eigenvalues converge to the true zero value while one of them remains positive. Additionally, the associated numerical eigenvector fields approximate the two harmonic fields of the Hodge Laplacian, one parallel to the $\theta$ direction and the other parallel to the $\phi$ direction. Before explaining the numerical observation for the positive eigenvalue, we first note that for a closed manifold, the discrete GMLS Laplace-Beltrami matrix always has a zero eigenvalue, with an associated constant eigenvector in the null space of the discrete matrix. However, this is not the case for a discrete vector Laplacian. A vector Laplacian may not have a zero eigenvalue (e.g., on the unit sphere), or if it does, the corresponding eigenvector field will not have constant components. Due to approximation errors, the leading numerical eigenvalue might be slightly greater than zero but it will converge to zero when a true zero eigenvalue exists, as shown in Fig. 4.
In the following PDE applications, we always test for the eigenvalue stability before solving the PDE. If a numerical eigenvalue $\hat{\lambda}_{0}$ has a small positive real part (as in the torus example), which characterizes the approximation error of the zero eigenvalue in this case, we will regularize the vector Laplacian matrix by subtracting a constant from its diagonal entries, $\mathbf{L}_{B}-\hat{\lambda}_{0}\mathbf{I}$, to guarantee the eigenvalue stability where $\mathbf{I}$ is an identity matrix.
5.2 Screened Poisson problems
In this subsection, we consider solving the following vector-valued screened Poisson problem for $\bm{u}\in\mathfrak{X}(M)$,
$(a-\Delta_{B})\bm{u}=\bm{f},\quad\mathbf{x}\in M,$ (49)
where $a=1>0$ and $\bm{f}\in\mathfrak{X}(M)$ are defined such that the problems are well-posed. For simplicity, we focus on the Bochner Laplacian in the following examples while Section 5.2.1 presents an example using the Hodge Laplacian. We note that the vector fields $\bm{f}$ and $\bm{u}$ must live in the tangent bundle $TM$ of the smooth manifold $M$. In the case of an unknown manifold, $\bm{f}$ is replaced by $\mathbf{\hat{T}}^{\top}\bm{F}$ in our numerical implementation where $\mathbf{\hat{T}}=[\bm{\hat{t}}_{1},\ldots,\bm{\hat{t}}_{d}]$ corresponds to the approximate tangent vectors and $\bm{F}\in\mathfrak{X}(\mathbb{R}^{n})$ is a vector field in ambient space. Thus, the vector field $\mathbf{\hat{T}}^{\top}\bm{F}$ and the approximate solution $\hat{\mathbf{u}}=(a-\mathbf{L}_{B})^{-1}(\mathbf{\hat{T}}^{\top}\bm{F})$ both live in the approximate tangent bundle $\widehat{TM}:=\cup_{\mathbf{x}\in M}\widehat{T_{\mathbf{x}}M}=\cup_{\mathbf{x}\in M}\mathrm{Span}\{\bm{\hat{t}}_{1}(\mathbf{x}),\ldots,\bm{\hat{t}}_{d}(\mathbf{x})\}$. To facilitate comparison with true solutions in the following numerical examples, we prescribe an analytic solution $\bm{u}$ and then approximate it, subjected to the manufactured $\bm{f}:=\mathbf{\hat{T}}^{\top}\bm{F}=\mathbf{\hat{T}}^{\top}(a-\bar{\Delta}_{B})\bm{U}$, where $(a-\bar{\Delta}_{B})\bm{U}$ is an extension of $(a-\Delta_{B})\bm{u}$, as defined in (38).
To verify convergence, we define the forward error (FE) and the inverse error (IE) in the ambient space $\mathbb{R}^{n}$ as
$\textbf{FE}=\max_{1\leq i\leq N}\|\hat{\textbf{T}}(\mathbf{x}_{i})(\mathbf{L}_{B}\mathbf{u}_{X_{M}})_{i}-\textbf{T}(\mathbf{x}_{i})\Delta_{B}\bm{u}(\mathbf{x}_{i})\|_{2},\mathrm{\ }\textbf{IE}=\max_{1\leq i\leq N}\|(\hat{\textbf{T}}\hat{\mathbf{u}})(\mathbf{x}_{i})-(\mathbf{T}\bm{u})(\mathbf{x}_{i})\|_{2},$
where $\|\cdot\|_{2}$ is the Euclidean distance, the true solution $\mathbf{u}_{X_{M}}=(\bm{u}(\mathbf{x}_{1}),\ldots,\bm{u}(\mathbf{x}_{N}))^{\top}\in\mathbb{R}^{dN\times 1}$ is a column vector, $(\mathbf{L}_{B}\mathbf{u}_{X_{M}})_{i}\in\mathbb{R}^{d\times 1}$ corresponds to the approximate value of $\Delta_{B}\bm{u}$ at $\mathbf{x}_{i}$, $\Delta_{B}\bm{u}(\mathbf{x}_{i})\in\mathbb{R}^{d\times 1}$ is the representation in the global basis at $\mathbf{x}_{i}$, $\hat{\mathbf{u}}$ is the approximate solution, $\mathbf{T}(\mathbf{x})$ $=[\bm{t}_{1}(\mathbf{x}),$$\ldots,\bm{t}_{d}(\mathbf{x})]$ $\in\mathbb{R}^{n\times d}$ and $\mathbf{\hat{T}}(\mathbf{x})=[\bm{\hat{t}}_{1}(\mathbf{x}),\ldots,\bm{\hat{t}}_{d}(\mathbf{x})]\in\mathbb{R}^{n\times d}$ correspond to the true and the approximate global bases, respectively. Here we push both tangent vectors on (approximate) manifolds forward to tangent vectors in $\mathbb{R}^{n}$ to make the FE and IE well-defined in the Euclidean norm.
5.2.1 Hodge Laplacian on 2D sphere in $\mathbb{R}^{3}$
In this numerical experiment, we solve the screened Poisson problem (49) on a 2D unit sphere associated with the Hodge Laplacian $\Delta_{H}$, which typically can be addressed using the FEEC framework. Here, we compare the performance of the GMLS approach proposed in this paper with the global RBF method [24] (a meshless approach) and the FEEC method [27] (a mesh-based approach), as discussed in Section 1. We set the true solution $\bm{u}=(x-x^{3})\mathbf{e}_{1}-(x^{2}y)\mathbf{e}_{2}-(x^{2}z)\mathbf{e}_{3}$ $\in\mathfrak{X}(M)$, where $\mathbf{x}=(x,y,z)$ is the Euclidean coordinate. For the GMLS approximation, we fix $K=50$ for $K$-nearest neighbors. In the RBF method, we select $\Psi(\mathbf{x})=\exp(-s^{2}\|\mathbf{x}-\mathbf{x}_{0}\|^{2})$ as the kernel function with a shape parameter $s=1$. For the surface FEM solution within the FEEC framework, we choose FEM spaces for differential forms as in [27], ensuring that the FEM spaces have the same polynomial degree as the surface approximation. Computations are performed using the NGSolve package [51]. For a consistent comparison, we generate triangle meshes over the surface with varying maximal mesh sizes. The mesh vertices are then extracted as a point cloud for the GMLS and RBF methods. The IEs are calculated as before at the vertices.
(a) (extr.) GMLS (b) (intr.) GMLS (c) FEEC &\&& RBF (d) time
Refer to caption Refer to caption Refer to caption Refer to caption
Figure 5: Screened Poisson problems on 2D sphere in $\mathbb{R}^{3}$ associated with the Hodge Laplacian. Random data with 6 trials. $K=50$ nearest neighbors are used. In panels (a) and (b), shown are the average IEs using extrinsic and intrinsic methods, respectively. In panels (c), shown are the IEs of the solutions using FEEC and RBF. In panels (d), shown are the comparison of the time cost among different methods. For time comparison, we use degree 2 for GMLS and FEEC.
In Figs. 5(a)-(c), we plot the $\mathbf{IE}$s as functions of $N$. The $\mathbf{IE}$s of the GMLS solutions decay at rates of $N^{-1}$ for $l=2,3$ and $N^{-2}$ for $l=4,5$, which agrees with the GMLS error estimates (7). In Fig. 5(c), we observe that for $l=2$ and $l=4$, the FEEC solution converges at the same rate as the GMLS methods but yields smaller errors. This behavior is expected as high-quality surface meshes can be easily constructed in this example. On the other hand, the RBF method provides the most accurate solution. In Fig. 5(d), we plot computational time as a function of $N$. We can see that the time cost of the RBF method increasing significantly faster than that of the GMLS and the FEEC methods. This is due to the fact that the RBF method produces a full matrix whereas the other two methods result in sparse matrices.
5.2.2 Bochner Laplacian on 2D torus in $\mathbb{R}^{9}$